# Learning with Incomplete Context: Linear Contextual Bandits with Pretrained Imputation

### v2

@@ E.1 Technical Lemmas
Appendix G Setup of the Lower Bound
@@
$\bm{Y}_{t}=\left(\bm{S}_{t}^{\top},W_{t}\right)^{\top}=\left(\bm{Q}^{\top}_{t},\bm{O}^{\top}_{t},W_{t}\right)^{\top}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}\times{{\mathbb{R}}}^{d_{\texttt{non}}}\times{{\mathbb{R}}}$ (G.1)
@@
$\bm{Y}_{t,-1}=\left(-\bm{Q}^{\top}_{t},\bm{0}^{\top}\right)^{\top}\in{{\mathbb{R}}}^{d_{0}},$ (G.2)
@@
$\mathbb{E}[W_{t}\mid\bm{S}_{t}]=f\left(\bm{O}_{t}\right)$ (G.3)
@@
$\mathcal{D}_{N}$ $=\left\{\left(\bm{S}^{(0)}_{i},W^{(0)}_{i}\right):i\in[N]\right\}$ (G.4)
@@
$\Theta:=\left\{\bm{\theta}\in{{\mathbb{R}}}^{d_{Y}}:\left\|\bm{\theta}\right\|_{2}\leq 1\right\}$ (G.5)
@@
$\Theta_{Q}:=\left\{\bm{\theta}_{Q}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}:\left\|\bm{\theta}_{Q}\right\|_{2}\leq\frac{\sqrt{3}}{2}\right\}.$ (G.6)
@@
$\bm{\theta}=\left(\bm{\theta}^{\top}_{Q},\bm{0}^{\top},\frac{1}{2}\right)^{\top}\in{{\mathbb{R}}}^{d_{Y}}\quad|\bm{\theta}_{Q,i}|=\sqrt{\frac{d_{\texttt{lin}}}{T}}.$ (G.7)
When $\bm{\theta}_{Q}\in\Theta_{Q}$, we have $\bm{\theta}\in\Theta$. Combining Equation (G.7) with the construction of $\bm{Y}_{t,1}$ in Equation (G.1) and $\bm{Y}_{t,-1}$ in Equation (G.2), we have
$\bm{\theta}^{\top}\bm{Y}_{t,1}=\bm{\theta}^{\top}_{Q}\bm{Q}_{t}+\frac{1}{2}W_{t},$ (G.8)
@@
Definition G.1 (Data Generating Process for Lower Bounds).
@@
$(V_{t},\bm{Q}_{t},\bm{O}_{t})=\begin{cases}\left(0,\ \bm{0},\ \bm{O}_{t}\right),&\text{with probability }\frac{1}{2}\\ \left(1,\ \bm{Q}_{t},\ \mathbf{o}_{0}\right),&\text{with probability }\frac{1}{2},\end{cases}$ (G.10)
where $\bm{O}_{t}\sim\mathbb{P}_{O}=\operatorname{Unif}([-1,1]^{d_{\texttt{non}}})$ and $\bm{Q}_{t}\sim\mathbb{P}_{Q}$ is a distribution specified in Equation (H.4) and $\mathbf{o}_{0}\in([-1,1]^{d_{\texttt{non}}})^{c}$ is an arbitrarily fixed vector such that $f(\mathbf{o}_{0})=0$ for $f$ given in Equation (G.3). Under the setup in Equation (G.10):
@@
    $\mathbb{P}_{f^{(a)}(\bm{O}_{t})}:=\mathbb{P}\left(R_{t}\mid A_{t}=a,\bm{S}_{t},V_{t}=0\right)$ (G.11)
@@
    $\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(a)}}:=\mathbb{P}\left(R_{t}\mid A_{t}=a,\bm{S}_{t},V_{t}=1\right)$ (G.12)
@@
Recall the historical data $\mathcal{D}_{N}$ defined in Equation (G.4). Let $\mathbb{P}_{f}$ denote the distribution of a pretraining sample $(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)},W_{i}^{(0)})$, with density
$p_{f}\left(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)},W_{i}^{(0)}\right)$ $=p_{f}\left(W^{(0)}_{i}\mid\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)}\right)p_{S}\left(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)}\right)$ (G.13)
@@
$p_{S}(\bm{Q}_{t},\bm{O}_{t})=\frac{1}{2}\delta_{0}(\bm{Q}_{t})p_{O}(\bm{O}_{t})+\frac{1}{2}p_{Q}(\bm{Q}_{t})\delta_{\mathbf{o}_{0}}(\bm{O}_{t}).$ (G.14)
@@
Assumption G.1.
For any $(\bm{\theta}_{1},f_{1}),(\bm{\theta}_{2},f_{2})\in\Theta\times\mathcal{F}_{\beta,L}$, the distributions in Equations (G.11) and (G.12) satisfy that
@@
Remark G.1.
Assumption G.1 can be satisfied by distributions such as Gaussian or Bernoulli.
@@
$\operatorname{KL}\left(\mathbb{P}^{(0)}_{f\left(\bm{O}_{i}^{(0)}\right)}\left(W^{(0)}_{i}\right)\|\mathbb{P}^{(0)}_{f^{\prime}\left(\bm{O}_{i}^{(0)}\right)}\left(W^{(0)}_{i}\right)\right)\leq C_{0}\left(f\left(\bm{O}_{i}^{(0)}\right)-f^{\prime}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}.$ (G.15)
@@
$p^{(t)}_{\bm{\theta},f,\pi}(\mathcal{D}_{N},\bm{Q}_{1},\bm{O}_{1},A_{1},R_{1},\cdots,\bm{Q}_{T},\bm{O}_{T},A_{T},R_{T})$ (G.16)
@@
where the equality follows from Equations (G.13) and (G.14). Additionally, let $\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}$ denote the expectation taken with respect to the joint density $p^{(t)}_{\bm{\theta},f,\pi}$.
We now state a formal definition of Theorem G.1.
Theorem G.1 (Formal Lower Bound).
Consider the data generating process given in Definition G.1. Suppose that $0<d_{\texttt{lin}}<c\sqrt{T}$ for some sufficiently small constant $c>0$ and $d_{\texttt{non}}>0$ is a constant. Fix a policy $\pi$. For any $(\bm{\theta},f)\in(\Theta,\mathcal{F}_{\beta,L})$, define
@@
where the joint density of the full observation history $\mathcal{H}_{t}$ up to round $t\in[T]$ is defined in Equation (G.16) and $R^{\star}_{t}=\max\left\{R(t,1),R(t,-1)\right\}$. For the class of functions $\mathcal{F}_{\beta,L}$ satisfies Assumptions 4.2, under Assumption G.1 and Equation (G.15), the expected cumulative regret is lower bounded by
$\sup_{\bm{\theta}\in\Theta,f\in\mathcal{F}_{\beta,L}}\mathcal{R}_{T}(\bm{\theta},f)=\Theta\left(TN^{-\frac{\beta}{2\beta+d_{\texttt{non}}}}\right)+\Theta\left(\sqrt{d_{\texttt{lin}}T}\right).$ (G.17)
Appendix H Proof of Theorem G.1

### v3

@@ G.1 Technical Lemmas
H.3 Robustness of Propositions 4.1 and 4.2 to Covariate Shift
The proofs above assume that the online contexts $\bm{S}_{t}$ follow the same marginal distribution as the historical data. Here we show that the rates for $\mathbb{E}[\sqrt{D_{t}}]$ in Propositions 4.1 and 4.2 are preserved when the marginal distribution of $\bm{S}_{t}$ shifts, provided the conditional mechanism $W_{t}\mid\bm{S}_{1:t}$ remains invariant and the marginal shift is mild.
Consider the linear model of Proposition 4.1 and suppose the online contexts are independent of the historical dataset $\mathcal{D}$ and satisfy $\bm{S}_{t}\sim\mathcal{N}(\bm{0},\bm{\Sigma}_{t})$ with a possibly time-varying covariance. The key step in the proof (see Equation (H.4)) involves bounding the distribution of the estimation error $\sum_{j=0}^{m}(\widehat{\bm{\beta}}_{j}-\bm{\beta}_{j})^{\top}\bm{S}_{t-j}$. Under the shifted marginal, this term is distributed as
$\sum_{j=0}^{m}(\widehat{\bm{\beta}}_{j}-\bm{\beta}_{j})^{\top}\bm{S}_{t-j}\;\sim\;\mathcal{N}\!\left(0,\;\sum_{j=0}^{m}(\widehat{\bm{\beta}}_{j}-\bm{\beta}_{j})^{\top}\bm{\Sigma}_{t-j}(\widehat{\bm{\beta}}_{j}-\bm{\beta}_{j})\right),$
and the same rates for $\mathbb{E}[\sqrt{D_{t}}]$ continue to hold as long as $\|\bm{\Sigma}_{t}\|_{\mathrm{op}}=O(1)$. Thus, even if the marginal distribution of $\bm{S}_{t}$ shifts in the online setting, the regret guarantees of Proposition 4.1 remain valid under uniformly bounded second moments. An analogous argument applies to Proposition 4.2: the nonparametric rate depends on the $L_{2}$ estimation error $\|\widehat{f}-f\|_{L_{2}}^{2}$, which is invariant to the online marginal as long as the conditional mechanism $W_{t}\mid\bm{S}_{t}$ is unchanged.
More broadly, the key condition is that the conditional mechanism $W_{t}\mid\bm{S}_{1:t}$ is invariant across the historical and online phases and the marginal shift in $\bm{S}_{1:t}$ is mild (e.g., bounded moments). Under these conditions, the rates in Section 4.3 remain unchanged. Investigating more general covariate-shift models, such as transfer-exponent conditions between historical and online distributions (Cai et al., 2024), is an interesting direction that requires substantially more technical care and is left for future work.
Appendix I Extension: A Confidence-Weighted Variant of PULSE-UCB
In the main text, we employ the standard OFUL algorithm in the online stage of PULSE-UCB as a simple, well-understood baseline. As demonstrated in Theorems 4.1 and 5.1, this plug-in approach already achieves near-optimal regret for our primary settings. However, since the context recovery step inherently introduces an imputation error $D_{t}$ for the feature at each step, one might ask whether we can make further improvements to mitigate this residual error.
To explore this, we draw inspiration from the literature on linear contextual bandits with corrupted rewards (He et al., 2022) to propose an alternative algorithmic variant. It is important to note that this variant is not a direct application of He et al. (2022); their work focuses on mitigating adversarial corruptions on the rewards, whereas our setting deals with misspecification in the contexts due to offline imputation. By adapting their confidence-weighted mechanism to our missing context framework, this alternative variant can potentially yield tighter regret bounds. Crucially, this improvement is conditional: it requires a reliable estimator for the total imputation error budget $\bar{C}$ a priori, and it is primarily beneficial when the imputation error $D_{t}$ varies substantially across time.
I.1 Algorithm Design: Robust PULSE-UCB
The primary modification in this robust variant lies in the construction of the confidence sets and the weighted regularized least-squares estimator. Specifically, we select actions via the modified confidence set $\widetilde{\texttt{BALL}}_{t}=\{\theta:\|\theta-\widetilde{\theta}_{t-1}\|_{\widetilde{\Sigma}_{t-1}}^{2}\leq\widetilde{\gamma}_{t-1}\}$, where the covariance matrix and the parameter estimator are updated using adaptive weights:
$\widetilde{\Sigma}_{t}=\lambda I+\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}\widehat{\phi}_{\tau,a_{\tau}}^{\top},\quad\widetilde{\theta}_{t}=\widetilde{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}R_{\tau}.$
The adaptive weight $w_{t}$ for each step is defined as:
$w_{t}=\min\left\{1,\frac{\alpha}{\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}}\right\},$
where $\alpha>0$ is a tuning parameter.
I.2 Theoretical Guarantees and Regret Comparison
We establish the theoretical guarantee for this robust variant as follows:
Proposition I.1 (Regret of Robust PULSE-UCB).
Let $\bar{C}$ be any known upper bound such that $\bar{C}\geq\sqrt{d}\sum_{t=1}^{T}\sqrt{D_{t}}$. By setting the tuning parameter $\alpha=\sqrt{d}/\bar{C}$ and the confidence radius $\widetilde{\gamma}_{t}=\gamma_{t}^{(0)}+3\alpha^{2}d\big(\sum_{t=1}^{T}\sqrt{D_{t}}\big)^{2}$, the cumulative regret of the robust variant satisfies, with high probability:
$R_{T}\leq R_{T}^{(lin)}+\widetilde{R}_{T}^{(imp)},$
where $R_{T}^{(lin)}$ matches the standard linear bandit regret term in Theorem 4.1, and the modified imputation penalty term is bounded by:
$\widetilde{R}_{T}^{(imp)}=12d\log^{3/2}\left(1+\frac{TB^{2}}{d\lambda}\right)\bar{C}.$
Proof.
Let $\phi_{t,a}$ denote the true feature vector and $\widehat{\phi}_{t,a}$ denote the imputed feature vector provided by the pretrained model. We can express the observed reward at round $t$ by explicitly isolating the stochastic noise and the context imputation error:
$R_{t}=\langle\theta^{\star},\phi_{t,a_{t}}\rangle+\eta_{t}=\langle\theta^{\star},\widehat{\phi}_{t,a_{t}}\rangle+c_{t}+\eta_{t},$
where the adversarial corruption term induced by imputation is defined as $c_{t}=\langle\theta^{\star},\phi_{t,a_{t}}-\widehat{\phi}_{t,a_{t}}\rangle$. Assuming the unknown environment parameter is bounded by $\|\theta^{\star}\|_{2}\leq S$ and the feature norm is bounded by $L$, the magnitude of this corruption is bounded by $|c_{t}|\leq S\|\phi_{t,a_{t}}-\widehat{\phi}_{t,a_{t}}\|_{2}\leq S\sqrt{D_{t}}$. Consequently, the total cumulative corruption over the horizon $T$ is $C=\sum_{t=1}^{T}|c_{t}|\leq S\sum_{t=1}^{T}\sqrt{D_{t}}$.
We first establish the concentration of the estimation error. The weighted ridge regression estimator $\widetilde{\theta}_{t}$ admits the closed-form representation:
$\widetilde{\theta}_{t}=\widetilde{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}(\langle\theta^{\star},\widehat{\phi}_{\tau,a_{\tau}}\rangle+c_{\tau}+\eta_{\tau})$
Subtracting $\theta^{\star}$ from both sides and multiplying by $\widetilde{\Sigma}_{t}$, the Mahalanobis norm of the estimation error satisfies:
$\|\widetilde{\theta}_{t}-\theta^{\star}\|_{\widetilde{\Sigma}_{t}}\leq\Big\|\widetilde{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}\eta_{\tau}\Big\|_{\widetilde{\Sigma}_{t}}+\Big\|\widetilde{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}c_{\tau}\Big\|_{\widetilde{\Sigma}_{t}}+\lambda\|\widetilde{\Sigma}_{t}^{-1}\theta^{\star}\|_{\widetilde{\Sigma}_{t}}.$
This decomposes the error into three absolute components: the stochastic error, the corruption error, and the regularization error. Because the adaptive weights $w_{\tau}$ are bounded by $1$, the stochastic error can be bounded using the standard self-normalized martingale concentration inequality by $R\sqrt{d\log((1+TL^{2}/\lambda)/\delta)}$ with probability at least $1-\delta$. The regularization error is strictly bounded by $\sqrt{\lambda}S$.
Crucially, we bound the corruption error utilizing the definition of the adaptive weight $w_{\tau}=\min\{1,\alpha/\|\widehat{\phi}_{\tau,a_{\tau}}\|_{\widetilde{\Sigma}_{\tau-1}^{-1}}\}$. By Cauchy-Schwarz and the weight constraint, we have:
$\Big\|\widetilde{\Sigma}_{t}^{-1}\sum_{\tau=1}^{t}w_{\tau}\widehat{\phi}_{\tau,a_{\tau}}c_{\tau}\Big\|_{\widetilde{\Sigma}_{t}}\leq\sum_{\tau=1}^{t}|c_{\tau}|w_{\tau}\|\widehat{\phi}_{\tau,a_{\tau}}\|_{\widetilde{\Sigma}_{t}^{-1}}\leq\sum_{\tau=1}^{t}|c_{\tau}|\alpha\leq\alpha C\leq\alpha S\sum_{t=1}^{T}\sqrt{D_{t}}.$
Therefore, with probability at least $1-\delta$, the true parameter $\theta^{\star}$ lies in the confidence set $\widetilde{\texttt{BALL}}_{t}$ for all $t\in[T]$, centered at $\widetilde{\theta}_{t-1}$ with the exact confidence radius:
$\beta_{T}=R\sqrt{d\log\left(\frac{1+TL^{2}/\lambda}{\delta}\right)}+\sqrt{\lambda}S+\alpha C.$
Conditioned on this high-probability event, we bound the instantaneous regret. Let $a_{t}^{\star}=\arg\max_{a\in\mathcal{A}}\langle\theta^{\star},\phi_{t,a}\rangle$. By the optimism of the action selection rule, $\langle\widetilde{\theta}_{t-1},\widehat{\phi}_{t,a_{t}^{\star}}\rangle+\beta_{T}\|\widehat{\phi}_{t,a_{t}^{\star}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}\leq\langle\widetilde{\theta}_{t-1},\widehat{\phi}_{t,a_{t}}\rangle+\beta_{T}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}$. The instantaneous regret satisfies:
$\text{reg}_{t}=\langle\theta^{\star},\phi_{t,a_{t}^{\star}}\rangle-\langle\theta^{\star},\phi_{t,a_{t}}\rangle\leq\langle\theta^{\star},\widehat{\phi}_{t,a_{t}^{\star}}\rangle-\langle\theta^{\star},\widehat{\phi}_{t,a_{t}}\rangle+2\max_{a\in\mathcal{A}}|c_{t,a}|.$
Applying the confidence bound and the optimism condition, this reduces to:
$\text{reg}_{t}\leq 2\beta_{T}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}+2\max_{a\in\mathcal{A}}|c_{t,a}|.$
Summing over the horizon $T$, the cumulative regret is strictly bounded by:
$R_{T}\leq\sum_{t=1}^{T}2\beta_{T}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}+2S\sum_{t=1}^{T}\sqrt{D_{t}}.$
To rigorously bound the sum of the weighted confidence widths, we partition the time indices into two disjoint subsets: $\mathcal{I}_{1}=\{t\in[T]:w_{t}=1\}$ and $\mathcal{I}_{2}=\{t\in[T]:w_{t}<1\}$. For the un-truncated subset $\mathcal{I}_{1}$, we directly apply the Cauchy-Schwarz inequality and the standard Elliptical Potential Lemma:
$\sum_{t\in\mathcal{I}_{1}}2\beta_{T}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}\leq 2\beta_{T}\sqrt{|\mathcal{I}_{1}|\sum_{t\in\mathcal{I}_{1}}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}^{2}}\leq 2\beta_{T}\sqrt{2dT\log\left(1+\frac{TL^{2}}{d\lambda}\right)}.$
This yields the standard linear bandit regret term $R_{T}^{(lin)}$.
For the truncated subset $\mathcal{I}_{2}$, the condition $w_{t}<1$ strictly implies $w_{t}=\alpha/\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}$. We algebraically rewrite the confidence width explicitly in terms of the weight:
$\sum_{t\in\mathcal{I}_{2}}2\beta_{T}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}=\sum_{t\in\mathcal{I}_{2}}2\beta_{T}\frac{w_{t}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}^{2}}{\alpha}.$
By applying the matrix determinant lemma and bounding the fractional terms, we obtain an absolute upper bound:
$\frac{2\beta_{T}}{\alpha}\sum_{t\in\mathcal{I}_{2}}w_{t}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}^{2}\leq\frac{2\beta_{T}}{\alpha}\sum_{t=1}^{T}\min\left(1,w_{t}\|\widehat{\phi}_{t,a_{t}}\|_{\widetilde{\Sigma}_{t-1}^{-1}}^{2}\right)\leq\frac{4\beta_{T}d}{\alpha}\log\left(1+\frac{TL^{2}}{d\lambda}\right).$
Finally, we substitute the specified tuning parameter $\alpha=\sqrt{d}/\bar{C}$, where the known upper bound satisfies $\bar{C}\geq\sqrt{d}\sum_{t=1}^{T}\sqrt{D_{t}}$. The corruption penalty embedded within the confidence radius satisfies $\alpha C\leq(\sqrt{d}/\bar{C})(S\bar{C}/\sqrt{d})=S$. Thus, the confidence radius simplifies to an inflation factor strictly independent of $T$: $\beta_{T}\leq R\sqrt{d\log((1+TL^{2}/\lambda)/\delta)}+\sqrt{\lambda}S+S$.
Substituting $\alpha$ into the upper bound for $\mathcal{I}_{2}$, the imputation-induced regret becomes:
$\widetilde{R}_{T}^{(imp)}=\frac{4\beta_{T}d\bar{C}}{\sqrt{d}}\log\left(1+\frac{TL^{2}}{d\lambda}\right)=4\sqrt{d}\beta_{T}\bar{C}\log\left(1+\frac{TL^{2}}{d\lambda}\right)+2S\frac{\bar{C}}{\sqrt{d}}.$
By expanding $\beta_{T}$ and grouping the leading polynomial and logarithmic terms, the penalty strictly respects the proposed theoretical bound $\widetilde{\mathcal{O}}(d\bar{C})$, verifying that the combination of bounded weights cleanly eliminates the $\sqrt{T}$ dependency strictly through the control of residual misspecification. This completes the proof. ∎
Discussion and Comparison: To understand the advantage of this robust variant, suppose the upper bound $\bar{C}$ is tightly chosen such that $\bar{C}=\Theta\big(\sqrt{d}\sum_{t=1}^{T}\sqrt{D_{t}}\big)$. In this case, the imputation-induced regret becomes $\widetilde{R}_{T}^{(imp)}=\widetilde{\mathcal{O}}\big(d^{3/2}\sum_{t=1}^{T}\sqrt{D_{t}}\big)$.
Recall that for the standard PULSE-UCB (Theorem 4.1), the corresponding imputation penalty is $R_{T}^{(imp)}=\widetilde{\mathcal{O}}\big(d^{3/2}\sqrt{T\sum_{t=1}^{T}D_{t}}\big)$. By the Cauchy–Schwarz inequality, we always have:
$\sum_{t=1}^{T}\sqrt{D_{t}}\leq\sqrt{T\sum_{t=1}^{T}D_{t}}.$
Consequently, the robust variant never exhibits a worse regret upper bound than the standard PULSE-UCB. More importantly, it can achieve strictly tighter bounds in environments where the imputation errors $D_{t}$ vary substantially (i.e., highly heterogeneous) over time.
Practical Trade-offs: While the robust variant provides tighter theoretical guarantees for heterogeneous errors, it requires the practitioner to specify the upper bound parameter $\bar{C}$ a priori. In contrast, if the errors $D_{t}$ are of the same order across time (as in the stationary settings discussed in Section 4.3), the Cauchy–Schwarz inequality is tight, and the regret rates of the two algorithms coincide. Therefore, practitioners can choose between the simpler standard PULSE-UCB and its robust variant based on whether a reliable estimate of $\bar{C}$ is available and the expected heterogeneity of the imputation errors.
Appendix J Setup of the Lower Bound
@@ I.2 Theoretical Guarantees and Regret Comparison
$\bm{Y}_{t}=\left(\bm{S}_{t}^{\top},W_{t}\right)^{\top}=\left(\bm{Q}^{\top}_{t},\bm{O}^{\top}_{t},W_{t}\right)^{\top}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}\times{{\mathbb{R}}}^{d_{\texttt{non}}}\times{{\mathbb{R}}}$ (J.1)
@@
$\bm{Y}_{t,-1}=\left(-\bm{Q}^{\top}_{t},\bm{0}^{\top}\right)^{\top}\in{{\mathbb{R}}}^{d_{0}},$ (J.2)
@@
$\mathbb{E}[W_{t}\mid\bm{S}_{t}]=f\left(\bm{O}_{t}\right)$ (J.3)
@@
$\mathcal{D}_{N}$ $=\left\{\left(\bm{S}^{(0)}_{i},W^{(0)}_{i}\right):i\in[N]\right\}$ (J.4)
@@
$\Theta:=\left\{\bm{\theta}\in{{\mathbb{R}}}^{d_{Y}}:\left\|\bm{\theta}\right\|_{2}\leq 1\right\}$ (J.5)
@@
$\Theta_{Q}:=\left\{\bm{\theta}_{Q}\in{{\mathbb{R}}}^{d_{\texttt{lin}}}:\left\|\bm{\theta}_{Q}\right\|_{2}\leq\frac{\sqrt{3}}{2}\right\}.$ (J.6)
@@
$\bm{\theta}=\left(\bm{\theta}^{\top}_{Q},\bm{0}^{\top},\frac{1}{2}\right)^{\top}\in{{\mathbb{R}}}^{d_{Y}}\quad|\bm{\theta}_{Q,i}|=\sqrt{\frac{d_{\texttt{lin}}}{T}}.$ (J.7)
When $\bm{\theta}_{Q}\in\Theta_{Q}$, we have $\bm{\theta}\in\Theta$. Combining Equation (J.7) with the construction of $\bm{Y}_{t,1}$ in Equation (J.1) and $\bm{Y}_{t,-1}$ in Equation (J.2), we have
$\bm{\theta}^{\top}\bm{Y}_{t,1}=\bm{\theta}^{\top}_{Q}\bm{Q}_{t}+\frac{1}{2}W_{t},$ (J.8)
@@
Definition J.1 (Data Generating Process for Lower Bounds).
@@
$(V_{t},\bm{Q}_{t},\bm{O}_{t})=\begin{cases}\left(0,\ \bm{0},\ \bm{O}_{t}\right),&\text{with probability }\frac{1}{2}\\ \left(1,\ \bm{Q}_{t},\ \mathbf{o}_{0}\right),&\text{with probability }\frac{1}{2},\end{cases}$ (J.10)
where $\bm{O}_{t}\sim\mathbb{P}_{O}=\operatorname{Unif}([-1,1]^{d_{\texttt{non}}})$ and $\bm{Q}_{t}\sim\mathbb{P}_{Q}$ is a distribution specified in Equation (K.4) and $\mathbf{o}_{0}\in([-1,1]^{d_{\texttt{non}}})^{c}$ is an arbitrarily fixed vector such that $f(\mathbf{o}_{0})=0$ for $f$ given in Equation (J.3). Under the setup in Equation (J.10):
@@
    $\mathbb{P}_{f^{(a)}(\bm{O}_{t})}:=\mathbb{P}\left(R_{t}\mid A_{t}=a,\bm{S}_{t},V_{t}=0\right)$ (J.11)
@@
    $\mathbb{P}_{\bm{\theta}_{Q}^{\top}\bm{Q}_{t}^{(a)}}:=\mathbb{P}\left(R_{t}\mid A_{t}=a,\bm{S}_{t},V_{t}=1\right)$ (J.12)
@@
Recall the historical data $\mathcal{D}_{N}$ defined in Equation (J.4). Let $\mathbb{P}_{f}$ denote the distribution of a pretraining sample $(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)},W_{i}^{(0)})$, with density
$p_{f}\left(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)},W_{i}^{(0)}\right)$ $=p_{f}\left(W^{(0)}_{i}\mid\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)}\right)p_{S}\left(\bm{Q}_{i}^{(0)},\bm{O}_{i}^{(0)}\right)$ (J.13)
@@
$p_{S}(\bm{Q}_{t},\bm{O}_{t})=\frac{1}{2}\delta_{0}(\bm{Q}_{t})p_{O}(\bm{O}_{t})+\frac{1}{2}p_{Q}(\bm{Q}_{t})\delta_{\mathbf{o}_{0}}(\bm{O}_{t}).$ (J.14)
@@
Assumption J.1.
For any $(\bm{\theta}_{1},f_{1}),(\bm{\theta}_{2},f_{2})\in\Theta\times\mathcal{F}_{\beta,L}$, the distributions in Equations (J.11) and (J.12) satisfy that
@@
Remark J.1.
Assumption J.1 can be satisfied by distributions such as Gaussian or Bernoulli.
@@
$\operatorname{KL}\left(\mathbb{P}^{(0)}_{f\left(\bm{O}_{i}^{(0)}\right)}\left(W^{(0)}_{i}\right)\|\mathbb{P}^{(0)}_{f^{\prime}\left(\bm{O}_{i}^{(0)}\right)}\left(W^{(0)}_{i}\right)\right)\leq C_{0}\left(f\left(\bm{O}_{i}^{(0)}\right)-f^{\prime}\left(\bm{O}_{i}^{(0)}\right)\right)^{2}.$ (J.15)
@@
$p^{(t)}_{\bm{\theta},f,\pi}(\mathcal{D}_{N},\bm{Q}_{1},\bm{O}_{1},A_{1},R_{1},\cdots,\bm{Q}_{T},\bm{O}_{T},A_{T},R_{T})$ (J.16)
@@
where the equality follows from Equations (J.13) and (J.14). Additionally, let $\mathbb{E}^{(t)}_{\bm{\theta},f,\pi}$ denote the expectation taken with respect to the joint density $p^{(t)}_{\bm{\theta},f,\pi}$.
We now state a formal definition of Theorem J.1.
Theorem J.1 (Formal Lower Bound).
Consider the data generating process given in Definition J.1. Suppose that $0<d_{\texttt{lin}}<c\sqrt{T}$ for some sufficiently small constant $c>0$ and $d_{\texttt{non}}>0$ is a constant. Fix a policy $\pi$. For any $(\bm{\theta},f)\in(\Theta,\mathcal{F}_{\beta,L})$, define
@@
where the joint density of the full observation history $\mathcal{H}_{t}$ up to round $t\in[T]$ is defined in Equation (J.16) and $R^{\star}_{t}=\max\left\{R(t,1),R(t,-1)\right\}$. For the class of functions $\mathcal{F}_{\beta,L}$ satisfies Assumptions 4.2, under Assumption J.1 and Equation (J.15), the expected cumulative regret is lower bounded by
$\sup_{\bm{\theta}\in\Theta,f\in\mathcal{F}_{\beta,L}}\mathcal{R}_{T}(\bm{\theta},f)=\Theta\left(TN^{-\frac{\beta}{2\beta+d_{\texttt{non}}}}\right)+\Theta\left(\sqrt{d_{\texttt{lin}}T}\right).$ (J.17)
Appendix K Proof of Theorem J.1